OpenStudy (anonymous):

What are the x– and y–intercepts of the graph of y = x2 – 10x + 21?

OpenStudy (raden):

to get x-intercept take y = 0 in other words x^2 – 10x + 21 = 0 solve for x, by factorization or use the quadratic formula

OpenStudy (raden):

to get y-intercept take x = 0 y = x^2 – 10x + 21 y = (0)^2 – 10(0) + 21 = ...

OpenStudy (raden):

dont forget they are represented as coordinat : for x-intercept : (...,0) for y-intercept : (0,...)

OpenStudy (anonymous):

I don't understand?

OpenStudy (raden):

x^2 – 10x + 21 = 0 this can be factored (x - 3)(x - 7) = 0 take each factor be equal zero, x-3 = 0 so, x=3 also x-7=0 x=7 thus, the x-intercept are (3,0) and (7,0)

OpenStudy (raden):

to get y-intercept take x = 0 y = x^2 – 10x + 21 y = (0)^2 – 10(0) + 21 = 0 - 0 + 21 = 21 thus, y-intercept at point (0,21)

OpenStudy (anonymous):

Thank you

OpenStudy (raden):

you're welcome